# Design of supercurrent diode by vortex phase texture

### v3

@@ Abstract
We investigate supercurrent nonreciprocal effects in a superconducting weak-link hosting distinct types of vortices. We demonstrate how the winding number of the vortex, its spatial configuration, and the shape of the superconducting lead can steer the sign and amplitude of the supercurrent rectification. We find a general criterion for the vortex pattern to maximize the rectification amplitude of the supercurrent. The underlying strategy is the search for specific vortex core positions yielding a vanishing amplitude of the supercurrent first harmonic. We also prove that supercurrent nonreciprocal effects can be used to diagnose high-winding vortex and to distinguish between different types of vorticity. Our results thus provide a toolkit to control the supercurrent rectification by means of vortex phase textures and nonreciprocal signatures to detect vortex states with nonstandard phase patterns.
I Introduction
@@
Vortices represent fundamental topological excitations in superfluid and superconductors. They have been predicted and successfully observed in a broad range of systems including superconductors [1], liquid helium [2, 3, 4], ultracold atomic gases [5], photon fields [6], and exciton-polariton condensates [7, 8]. Vortices are generally characterized by a quantized phase-winding and a suppressed order parameter at their core. A gradient of the phase $\phi$ of the superconducting order parameter $\Delta=|\Delta|\exp{(i\phi)}$ yields a circulating supercurrent around the core of the vortex whereas the amplitude is vanishing, i.e. $|\Delta|\rightarrow 0$. For conventional superconductors, due to the single-valuedness of the superconducting order parameter, the winding number associated with the phase gradient is forced to be an integer $V_{0}$. In principle, $V_{0}$ can assume values larger than one. Apart from having $V_{0}$ vortices with winding number equal to one, a giant vortex with winding number $V_{0}$ can also realize. A giant vortex is expected to be relevant in small superconductors with confined geometries. To date, probing of such unconventional vortex state has been mostly addressed by magnetic means in suitably tailored geometric configurations, e.g. by Hall probe [9, 10, 11] and scanning SQUID microscopy [12], or by tunneling microscopy and spectroscopy [13, 14, 15, 16, 17].
@@
Here, we unveil a specific relation between the occurrence of vortex states with any given winding number in a Josephson weak-link and the rectification of the supercurrent flowing across the junction. Supercurrent rectification is a timely problem at the center of intense investigation [18, 19, 20, 21, 22, 23, 24]. A large body of work devoted to supercurrent rectification focuses on superconducting states marked by linear phase gradients with, for instance, Cooper pair momentum [25, 26, 27], spin-flipper by ferromagnets [28], or helical phases [29, 30, 31, 32, 33], as well as screening currents [34, 35], and supercurrent related to self-field [36, 37] or back-action mechanisms [38]. Vortices or circular phase gradients, associated to conventional winding, are also expected to yield supercurrent diode effects, due to the induced Josephson phase shift [39], and their role has been investigated for a variety of physical configurations [37, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. However, whether and how superconducting phase patterns with nontrivial winding for the vorticity can be probed by nonreciprocal response, are problems not yet fully uncovered.
@@
The structure of the paper is as follows. Sect. II provides the details of the model Hamiltonian, the formulation and the methodology employed in this study. In Sect. III we present the results for the supercurrent rectification in the case of one and two vortices within the superconducting leads for different values of the vortex winding. Sect. III is devoted to the analysis of the supercurrent rectification assuming that the supercurrent is vanishing at the boundary of the junction. In Section IV, we examine how the boundary affects the vortex phase texture and assess the resulting behavior of supercurrent rectification. The conclusions are reported in Sect. V.
II Model and methodology
@@
Figure 1: (a) Schematic of a Josephson junction in the lattice model with a vortex phase texture. (b) A representative nonreciprocal current phase relation. (c)(d) Sketch of the superconducting weak link with a vortex placed on the left side of the junction. We show two representative vortex positions along the lateral direction. Since the supercurrent pattern is spatially modified by the phase vortex, it can become nonreciprocal, i.e. the forward supercurrent $I_{+}$ is different from the backward one $I_{-}$. The vortex winding, $V^{\mathrm{L}}_{0}$, can take any integer number $l$.
II.1 Model Hamiltonian
We consider a planar superconducting weak-link with the geometry shown in Fig. 1 (a). For convenience, the system size is expressed as $N_{y}a\times(N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x})a$ with $N_{y}$ being the lateral number of sites, $N^{\mathrm{L,R,N}}_{x}$ the site numbers in the left ($L$), right ($R$) and normal ($N$) region of the superconducting weak-link, and $a$ is the lattice length, respectively. The superconducting order parameter on the left and right side of the junction is given by $\Delta_{\mathrm{L,R}}e^{i\varphi_{\mathrm{L,R}}}$.
@@ II.1 Model Hamiltonian
^$\hat{\mathcal{H}}=\sum_{j_{x},j_{y}}\sum_{j^{\prime}_{x},j^{\prime}_{y}}\hat{C}^{\dagger}(j_{x},j_{y})\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)\hat{C}(j^{\prime}_{x},j^{\prime}_{y}),$ (1)
@@
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}+$
@@
$\hat{H}^{\mathrm{R}}$ $=\sum^{N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}}_{j_{x}=-N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+1}\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[(-\varepsilon)c^{\dagger}_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}+$
@@
with the on-site energy term $\varepsilon=-0.25t$, the energy gap function with spin-singlet $s$-wave state $\Delta_{\mathrm{L,R}}e^{\varphi_{\mathrm{L,R}}}$, and the nearest neighbor hopping terms $t_{x,y}$. In the present study, the pairing amplitude is on-site and included in the local term of Hamiltonian for the left and right-side superconductors. The tunneling Hamiltonian is expressed as
@@
$\hat{H}^{\mathrm{R}}_{J}$ $=\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[-t_{\mathrm{int}}tc^{\dagger}_{-L^{\mathrm{L}}_{x}+L^{\mathrm{N}}_{x}+1,j_{y},\uparrow}c_{-L^{\mathrm{L}}_{x}+L^{\mathrm{N}}_{x},j_{y},\uparrow}]$
@@
with $t_{\mathrm{int}}$ the charge transfer amplitude at the interface setting out the transparency of the junction.
II.2 Josephson current and rectification
Next, we consider how the Josephson current flowing in the junction [Fig. 1 (a)] is evaluated. The current phase relation of the Josephson current is obtained by evaluating the variation of the free energy $F$ with respect to the phase bias across the junction:
@@ II.2 Josephson current and rectification
with the free energy in a Josephson junction at zero temperature evaluated as
@@
Here, $N_{x}=N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}$ and $N_{y}$ denote the total number of sites along the $x$-direction and along the $y$-direction, and $E(\varphi)$ and $|\Phi\rangle$ stand for the eigenvalue and eigenstate of $\hat{H}(x,y;x^{\prime},y^{\prime};\varphi)$. We numerically obtain $E(\varphi)$ by the full diagonalization of $\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)$ in Eq. (2). The computational analysis is performed at zero temperature, but a thermal change does not alter qualitatively the results. For our purposes, we recall that if time-reversal is broken then the supercurrent can be expanded in even and odd harmonics with respect to the phase bias variable as
@@
with $I_{+(-)}$ the maximum amplitude of the supercurrent for forward (backward) directions, respectively. For the presented results, we set the energy gap amplitude as $|\Delta_{0}|=0.02t$ ($t$ is the electron hopping amplitude), the maximum Josephson current value, without vortex, as $I_{\mathrm{0}}=0.012|\Delta_{0}|$, and the size of the superconducting leads as $N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=\alpha N_{y}$ with $\alpha$ being the aspect ratio setting out a square or rectangular shape of the superconductors in the junction. The computation is performed for $N_{y}=30$, and $z^{\mathrm{L}}_{0}=10a$ for the size of the vortex core. A variation of these lengths does not affect the results, thus, for clarity, we introduce the coordinates $(j_{x},j_{y})$ in the Josephson junction [Fig. 2 (a)] (see details in Appendix A).
@@
Figure 2: (a) Schematic illustration of the Josephson junction with real space coordinates. The analysis is performed for a vortex configuration having winding number $V^{\mathrm{L}}_{0}=+l$ with $l=1,2,3$. Evolution of the rectification amplitude $\eta$ with regard to the vortex core coordinates for different winding numbers: (b,c) $V^{\mathrm{L}}_{0}=1$, (f,g) $V^{\mathrm{L}}_{0}=2$, and (j,k) $V^{\mathrm{L}}_{0}=3$. Spatially resolved harmonics of the supercurrent: $I_{1}$ (red-solid line), $I_{2}$ (green-dotted), and $J_{1}$ (blue-dashed) indicate the odd-parity first harmonic, the odd-parity second harmonic, and the even-parity first harmonic amplitude, respectively. Longitudinal scan: $I_{1}$, $I_{2}$, and $J_{1}$ at a given $y^{L}_{0}$ as a function of $x^{L}_{0}$ for (d) $V^{\mathrm{L}}_{0}=1$, (h) $V^{\mathrm{L}}_{0}=2$, and (l) $V^{\mathrm{L}}_{0}=3$. Lateral scan: $I_{1}$, $I_{2}$, and $J_{1}$ at a given $x^{L}_{0}$ as a function of $y^{L}_{0}$ for (e) $V^{\mathrm{L}}_{0}=1$, (i) $V^{\mathrm{L}}_{0}=2$, and (m) $V^{\mathrm{L}}_{0}=3$. The gray dotted lines refer to the position of the maximal rectification and are a guide to indicate the values $I_{1}$, $I_{2}$, and $J_{1}$. We set $y^{\mathrm{L}}_{0}$ as (b)(d) $0.5a$ and (f)(h)(j)(l) $5.5a$, and $x^{\mathrm{L}}_{0}$ as (c)(e)(g)(i) $-23a$ and (k)(m) $-6a$. The aspect ratio is (b)-(i) $\alpha=3/2$ and (j)-(m) $\alpha=1$. The maximal rectification $\eta$ occurs for vortex core positions corresponding to a supercurrent with $I_{1}$, $I_{2}$, and $J_{1}$ components that are comparable in size. The sign change of $\eta$ is related to the vanishing of $J_{1}$ and to the zeros of $I_{1}$ when the amplitude is comparable to $J_{1}$. Multiple sign reversals of $\eta$ are observed for $V^{\mathrm{L}}_{0}=3$.
@@
Figure 3: Amplitude of the first harmonics of the Josephson current versus the vortex core positions for different vortex winding: (a,b) $V^{\mathrm{L}}_{0}=1$, (c,d) $V^{\mathrm{L}}_{0}=2$, (e,f) $V^{\mathrm{L}}_{0}=3$. $\bar{I}_{1}$ and $\bar{J}_{1}$ indicate the odd- and even-parity first harmonics of the supercurrent evaluated by the direct Cooper pairs tunneling. Gray-dotted and dashed lines stand for the vortex core positions, $y^{\mathrm{L}}_{0}$ and $x^{\mathrm{L}}_{0}$, evaluated in Fig. 2. The aspect ratio is (a)-(d) $\alpha=3/2$ and (e,f) $\alpha=1$.
III Supercurrent Rectification: superconducting lead larger than the vortex phase coherence
In this section we study the supercurrent rectification by examining one and two vortices situated within the superconducting lead of the junction and assuming that it can be marked by different winding numbers. Here, we will first consider the case where the size of the vortex is much smaller than the superconducting lateral dimension so that the edges do not affect the vortex phase texture. Then, in Sect. IV we also analyze the rectification for the case of small size superconductors by imposing that the supercurrent is vanishing at the edge.
III.1 Single vortex configuration
@@
To investigate the rectification properties of the supercurrent in the junction, we start by considering a single vortex configuration assuming a variable winding number and core position as shown in Figs. 1 (b,c). The vortex is placed on the left side of the junction but the results for vortices on the right side of the junction can be simply obtained using that inversion symmetry leads to a sign variation of the rectification amplitude (see Appendix A). Antivortex configurations through mirror and time-reversal transformations can also be directly deduced from the results of the single vortex (see Appendix A). To simulate the vortex state with the core at a given position, $\bm{r}^{\mathrm{L}}_{0}=(x_{0}^{\mathrm{L}},y_{0}^{\mathrm{L}})$, the phase value at the site $\bm{j}$ in the left-side superconductor is given by
$\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})=V^{\mathrm{L}}_{0}\arg[(j_{x}a-x_{0}^{\mathrm{L}}a)+i(j_{y}a-y_{0}^{\mathrm{L}})],$ (13)
@@
and $\Delta_{\mathrm{R}}=|\Delta_{0}|$ with $|\Delta_{0}|$ being the superconducting energy gap amplitude. In the presence of a phase vortex texture, the amplitude of the pair potential $\Delta_{\mathrm{L}}$ is modified by $\tilde{\Theta}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ for each site $\bm{j}$:

### v5

# Supercurrent diode with high winding vortex

@@ Abstract
Nonreciprocal supercurrent refers to the phenomenon where the maximum dissipationless current in a superconductor depends on its direction of flow. This asymmetry underlies the operation of superconducting diodes and is often associated with the presence of vortices. Here, we investigate supercurrent nonreciprocal effects in a superconducting weak-link hosting distinct types of vortices. We demonstrate how the winding number of the vortex, its spatial configuration, and the shape of the superconducting lead can steer the sign and amplitude of the supercurrent rectification. We identify a general criterion for optimizing the rectification amplitude based on vortex patterns, focusing on configurations where the first harmonic of the supercurrent vanishes. We prove that supercurrent nonreciprocal effects can be used to diagnose high-winding vortex and to distinguish between different types of vorticity. Our results provide a toolkit for controlling supercurrent rectification through vortex phase textures and detecting unconventional vortex states.
Introduction
@@
Vortices represent fundamental topological excitations in superfluids and superconductors. They have been predicted and successfully observed in a broad range of systems, including superconductors [1], liquid helium [2, 3, 4], ultracold atomic gases [5], photon fields [6], and exciton-polariton condensates [7, 8]. Vortices are generally characterized by a quantized phase winding and a suppressed order parameter at their core. A gradient of the phase $\phi$ of the superconducting order parameter $\Delta=|\Delta|\exp(i\phi)$ yields a circulating supercurrent around the core of the vortex, whereas the amplitude is vanishing, i.e. $|\Delta|\rightarrow 0$. For conventional superconductors, due to the single-valuedness of the superconducting order parameter, the winding number associated with the phase gradient is forced to be an integer $V_{0}$. In principle, $V_{0}$ can assume values larger than one. Apart from having $V_{0}$ vortices with winding number equal to one, a giant vortex with winding number $V_{0}$ can also be realized. A giant vortex is expected to be relevant in small superconductors with confined geometries. To date, probing of such unconventional vortex state has been mostly addressed by magnetic means in suitably tailored geometric configurations, e.g. by Hall probe [9, 10, 11] and scanning SQUID microscopy [12], or by tunneling microscopy and spectroscopy [13, 14, 15, 16, 17].
@@
Here, we unveil a specific relation between the occurrence of vortex states with any given winding number in a Josephson weak-link and the rectification of the supercurrent flowing across the junction. Supercurrent rectification is a timely problem at the center of intense investigation [18, 19, 20, 21, 22, 23, 24]. A large body of work devoted to supercurrent rectification focuses on superconducting states marked by linear phase gradients with, for instance, Cooper pair momentum [25, 26, 27], spin-flipper by ferromagnets [28], or helical phases [29, 30, 31, 32, 33], as well as screening currents [34, 35], and supercurrent related to self-field [36, 37] or back-action mechanisms [38]. Vortices or circular phase gradients, associated to conventional winding, are also expected to yield supercurrent diode effects, due to the induced Josephson phase shift [39], and their role has been investigated for a variety of physical configurations [37, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. However, whether and how superconducting phase patterns with nontrivial winding for the vorticity can be probed by nonreciprocal response, which are problems not yet fully uncovered.
@@
Model and methodology
@@
Figure 1: Image of Josephson junctions with a vortex. (a) Schematic of a Josephson junction in the lattice model with a vortex phase texture. SC-L, N, and SC-R mean the Left-side superconductor (SC), the Normal metal, and the Right-side SC. $N^{\mathrm{L}}_{x}$, $N^{\mathrm{R}}_{x}$, $N^{\mathrm{N}}_{x}$, $N_{y}$, and $a$ denote the number of sites in the left and right-side SCs, and in the normal metal along the $x$-direction, and number of sites along the $y$-direction, and lattice constant. $\bm{r}^{\mathrm{L}}_{0}=(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})$ is the core position. (b) A representative nonreciprocal current phase relation with $\varphi$ being the phase bias across the junction. (c)(d) Sketch of the superconducting weak link with a vortex placed on the left side of the junction. We show two representative vortex positions along the lateral direction. Since the supercurrent pattern is spatially modified by the phase vortex, it can become nonreciprocal, i.e. the forward supercurrent $I_{+}$ is different from the backward one $I_{-}$. The vortex winding, $V^{\mathrm{L}}_{0}$, can take any integer number $l$.
Model Hamiltonian. We consider a planar superconducting weak-link with the geometry shown in Fig. 1 (a). For convenience, the system size is expressed as $N_{y}a\times(N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x})a$ with $N_{y}$ being the lateral number of sites, $N^{\mathrm{L,R,N}}_{x}$ the site numbers in the left ($L$), right ($R$) and normal ($N$) region of the superconducting weak-link, and $a$ is the lattice length, respectively. The superconducting order parameter on the left and right side of the junction is given by $\Delta_{\mathrm{L,R}}e^{i\varphi_{\mathrm{L,R}}}$.
@@
$\hat{\mathcal{H}}=\sum_{j_{x},j_{y}}\sum_{j^{\prime}_{x},j^{\prime}_{y}}\hat{C}^{\dagger}(j_{x},j_{y})\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)\hat{C}(j^{\prime}_{x},j^{\prime}_{y}),$ (1)
@@
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
@@
$\hat{H}^{\mathrm{R}}$ $=\sum^{N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}}_{j_{x}=N^{\mathrm{N}}_{x}+1}\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[(-\varepsilon)c^{\dagger}_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
@@
with the on-site energy term $\varepsilon=-0.25t$, the energy gap function with spin-singlet $s$-wave state $\Delta_{\mathrm{L,R}}e^{i\varphi_{\mathrm{L,R}}}$, and the nearest neighbor hopping terms $t_{x,y}$. In the present study, the pairing amplitude is on-site and included in the local term of the Hamiltonian for the left and right-side superconductors. We point out that the position dependence of the order parameter $\Delta_{\mathrm{R}}(j_{x},j_{y};\varphi)$ takes into account the change in the phase and amplitude due to the presence of the vortices. This is a convenient approach that can be generally applied for analyzing the effects of nonstandard vortices without explicitly incorporating the source responsible for generating the vortex. The tunneling Hamiltonian is expressed as
@@
$\hat{H}^{\mathrm{R}}_{J}$ $=\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[-t_{\mathrm{int}}tc^{\dagger}_{N^{\mathrm{N}}_{x}+1,j_{y},\uparrow}c_{N^{\mathrm{N}}_{x},j_{y},\uparrow}]$
@@
with $t_{\mathrm{int}}=0.90$ the charge transfer amplitude at the interface, setting out the transparency of the junction.
Josephson current and rectification. Next, we consider how the Josephson current flowing in the junction [Fig. 1 (a)] is evaluated. The current phase relation of the Josephson current is obtained by evaluating the variation of the free energy $F$ with respect to the phase bias across the junction, i.e. $\varphi$:
@@
with the free energy in a Josephson junction at zero temperature, evaluated as
@@
Here, $N_{x}=N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}$ and $N_{y}$ denote the total number of sites along the $x$-direction and along the $y$-direction, and $E(\varphi)$ and $|\Phi\rangle$ stand for the eigenvalue and eigenstate of $\hat{H}(x,y;x^{\prime},y^{\prime};\varphi)$. We numerically obtain $E(\varphi)$ by the full diagonalization of $\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)$ in Eq. (2). The computational analysis is performed at zero temperature, but a thermal change does not qualitatively alter the results. For our purposes, we recall that if time-reversal is broken, then the supercurrent can be expanded in even and odd harmonics with respect to the phase bias variable as
@@
with $I_{+(-)}$ the maximum amplitude of the supercurrent for forward (backward) directions, respectively. For the presented results, we set the energy gap amplitude as $|\Delta_{0}|=0.02t$ ($t$ is the electron hopping amplitude), the maximum Josephson current value, without vortex, as $I_{\mathrm{0}}=0.012|\Delta_{0}|(\frac{2e}{\hbar})$, and the size of the superconducting leads as $N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=\alpha N_{y}$ with $\alpha$ being the aspect ratio setting out a square or rectangular shape of the superconductors in the junction. The computation is performed for $N_{y}=30$, and $z^{\mathrm{L}}_{0}=10a$ for the size of the vortex core. A variation of these lengths does not affect the results; thus, for clarity, we introduce the coordinates $(j_{x},j_{y})$ in the Josephson junction [Fig. 2 (a)].
@@
Figure 2: Space dependence of rectification amplitudes with a vortex. (a) Schematic illustration of the Josephson junction with real space coordinates. The analysis is performed for a vortex configuration having winding number $V^{\mathrm{L}}_{0}=+l$ with $l=1,2,3$. SC-L, N, and SC-R mean the Left-side superconductor (SC), the Normal metal, and the Right-side SC. $N^{\mathrm{L}}_{x}$, $N^{\mathrm{R}}_{x}$, $N^{\mathrm{N}}_{x}$, $N_{y}$, and $a$ denote the number of sites in the left and right-side SCs, and in the normal metal along the $x$-direction, and number of sites along the $y$-direction, and lattice constant. Evolution of the rectification amplitude $\eta$ with regard to the vortex core coordinates for different winding numbers: (b,c) $V^{\mathrm{L}}_{0}=1$, (f,g) $V^{\mathrm{L}}_{0}=2$, and (j,k) $V^{\mathrm{L}}_{0}=3$. Spatially resolved harmonics of the supercurrent: $I_{1}$ (red-solid line), $I_{2}$ (green-dotted), and $J_{1}$ (blue-dashed) indicate the odd-parity first harmonic, the odd-parity second harmonic, and the even-parity first harmonic amplitude, respectively. Longitudinal scan: $I_{1}$, $I_{2}$, and $J_{1}$ at a given $y^{\mathrm{L}}_{0}$ as a function of $x^{\mathrm{L}}_{0}$ for (d) $V^{\mathrm{L}}_{0}=1$, (h) $V^{\mathrm{L}}_{0}=2$, and (l) $V^{\mathrm{L}}_{0}=3$. Lateral scan: $I_{1}$, $I_{2}$, and $J_{1}$ at a given $x^{\mathrm{L}}_{0}$ as a function of $y^{\mathrm{L}}_{0}$ for (e) $V^{\mathrm{L}}_{0}=1$, (i) $V^{\mathrm{L}}_{0}=2$, and (m) $V^{\mathrm{L}}_{0}=3$. The gray dotted lines refer to the position of the maximal rectification and are a guide to indicate the values $I_{1}$, $I_{2}$, and $J_{1}$. We set $y^{\mathrm{L}}_{0}$ as (b)(d) $0.5a$ and (f)(h)(j)(l) $5.5a$, and $x^{\mathrm{L}}_{0}$ as (c)(e)(g)(i) 23a-23a and (k)(m) 6a-6a. The aspect ratio is (b)-(i) $\alpha=3/2$ and (j)-(m) $\alpha=1$. The maximal rectification $\eta$ occurs for vortex core positions corresponding to a supercurrent with $I_{1}$, $I_{2}$, and $J_{1}$ components that are comparable in size. The sign change of $\eta$ is related to the vanishing of $J_{1}$ and to the zeros of $I_{1}$ when the amplitude is comparable to $J_{1}$. Multiple sign reversals of $\eta$ are observed for $V^{\mathrm{L}}_{0}=3$. Parameters: $|\Delta_{0}|=0.02t$ (superconducting energy gap amplitude), $t_{\mathrm{int}}=0.90$ (transparency at the interface), $N^{\mathrm{N}}_{x}=10$, $N_{y}=30$, and $z^{\mathrm{L}}_{0}=10a$ (vortex size).
@@
Figure 3: Evaluation of the first and second harmonics and the rectifications as a function of the space with a vortex. Amplitude of the first harmonics of the Josephson current and the rectification versus the vortex core positions $\bm{r}^{\mathrm{L}}_{0}=(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})$ for different vortex winding $V^{\mathrm{L}}_{0}$: (a,b,c) $V^{\mathrm{L}}_{0}=1$, (d,e,f) $V^{\mathrm{L}}_{0}=2$, (g,h,i) $V^{\mathrm{L}}_{0}=3$. $\bar{I}_{1}$, $\bar{J}_{1}$, and $\eta$ indicate the odd- and even-parity first harmonics of the supercurrent evaluated by the direct Cooper pairs tunneling and the amplitude of the rectification. The color bars indicate the amplitude of (a,d,g) $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$, (b,e,h) $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$, and (c,f,i) the rectification amplitude $\eta$. In (a,b,d,e,g,h), gray-dotted and dashed lines stand for the vortex core positions, $y^{\mathrm{L}}_{0}$ and $x^{\mathrm{L}}_{0}$, evaluated in Fig. 2. The aspect ratio is (a)-(f) $\alpha=3/2$ and (g)-(i) $\alpha=1$. In (c)(f)(i), the rectification amplitude is evaluated by scanning all the positions of the vortex cores by performing the computation of the supercurrent for the weak link assuming the following parameters: $|\Delta_{0}|=0.02t$ (superconducting energy gap amplitude), $t_{\mathrm{int}}=0.90$ (transparency at the interface), $N^{\mathrm{N}}_{x}=10$, $N_{y}=30$, and $z^{\mathrm{L}}_{0}=10a$ (vortex size).
Supercurrent Rectification
In this section, we study the supercurrent rectification by examining one and two vortices situated within the superconducting lead of the junction and assuming that the core position can be varied and it can be marked by different winding numbers. We would like to point out that to realize nonreciprocal superconducting phenomena, it is essential to break both inversion and time-reversal symmetries. This condition is crucial for establishing nonreciprocal behavior in superconducting transport, and it can be directly inferred in a superconducting weak link, for example, from the parity properties of the current-phase relation with respect to the phase bias. In the examined system, these symmetries are broken by the presence of vortices in the superconducting leads. Specifically, when analyzing the junction with either one or two vortices, the system lacks a center of inversion and is not invariant under time-reversal symmetry. For example, considering a single vortex on one side of the junction, there are no inversion centers, and the phase pattern does not remain invariant under time reversal. These symmetry considerations directly influence the current-phase relation of the Josephson junction. Specifically, the relation features a first harmonic with even parity, which implicitly reflects the broken time-reversal symmetry. Additionally, the second harmonic contains both even- and odd-parity components with respect to the applied phase bias. Although the odd-parity component can exist regardless of the vortex presence when the junction is outside the tunneling regime, the even-parity contribution arises from the vortex-induced symmetry breaking. Then, we note that the rectification amplitude is generally nonzero when a nontrivial even-parity first harmonic coexists with the second harmonic component. Our results thus will demonstrate that the vortex plays a role in the rectification process of the supercurrent by generating a nonzero even-parity first harmonic term. In particular, the position of the vortex affects both the amplitude of the first harmonic term and the second harmonic component, too, thus resulting in a modulation of the sign and amplitude of the supercurrent rectification. Here, we adopt a superconducting phase profile designed to qualitatively capture the main features of the giant vortex as observed in relevant experiments [51, 52, 13, 53, 15] and discussed in several theoretical studies [54, 55, 56, 57, 58, 59, 60]. However, this approach does not incorporate the boundary-induced corrections required to fully satisfy supercurrent conservation at the sample edges. While such corrections are known to play a crucial role in stabilizing high-winding-number vortices, particularly in finite-size systems [55, 56, 57, 58, 59, 60] or in the presence of strong pinning [54], we point out that a self-consistent treatment of the superconducting phase, accounting for boundary effects, lies beyond the scope of the present study.
@@
Single vortex configuration. To investigate the rectification properties of the supercurrent in the junction, we start by considering a single vortex configuration assuming a variable winding number and core position as illustrated in Fig. 1. The vortex is placed on the left side of the junction, but the results for vortices on the right side of the junction can be directly obtained by applying an inversion symmetry transformation and considering that it leads to a sign variation of the rectification amplitude (see Supplementary Info - Section B). Antivortex configurations through mirror and time-reversal transformations can also be directly deduced from the results of the single vortex (see Supplementary Info - Section B). To simulate the vortex state with the core at a given position, $\bm{r}^{\mathrm{L}}_{0}=(x_{0}^{\mathrm{L}},y_{0}^{\mathrm{L}})$, the phase value at the site $\bm{j}$ in the left-side superconductor is given by (Fig.  2 (a))
$\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})=V^{\mathrm{L}}_{0}\arg[(j_{x}a-x_{0}^{\mathrm{L}})+i(j_{y}a-y_{0}^{\mathrm{L}})],$ (13)
@@
and $\Delta_{\mathrm{R}}=|\Delta_{0}|$ with $|\Delta_{0}|=0.02t$ being the superconducting energy gap amplitude. In the presence of a phase vortex texture, the amplitude of the pair potential $\Delta_{\mathrm{L}}$ is modified by $\tilde{\Theta}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ for each site $\bm{j}$: